# A Lightweight Method for Defending Against UAF Vulnerabilities

### v1

@@ 4.1.1. Program Abstraction
Table 1. Analysis Domains of Our Pointer Analysis.
@@
$m$in $\mathcal{M}$subseteq $\mathcal{G}$ program functions
@@
The domains related to our pointer analysis are shown in Table 1, mainly including variables, functions, types, and objects. Variables in the program $\mathcal{V}$ can be divided into two types: address-taken variables $\mathcal{A}$, which can be accessed through pointers, and top-level variables $\mathcal{P}$, which are never accessed through pointers. Address-taken variables $\mathcal{A}$ include abstract objects $\mathcal{O}$ and fields of abstract objects $\mathcal{Y}$. Heap objects are a type of address-taken object, and our method for abstracting heap objects is to treat each heap allocation point as a distinct object. We represent object fields by creating a sub-object for each field and establishing a relationship between the sub-object and its parent object. The creation of sub-objects is based on the offset of their fields relative to the base object.
Table 2. LLVM-like Instructions.
@@
Our pointer analysis is based on the LLVM IR of programs. LLVM defines a rich set of instructions, but only a small subset is relevant to our pointer analysis. Table 2 shows the set of instructions related to our pointer analysis. STACK/GLOBAL instructions allocate stack or global objects. HEAP instructions are used for heap object allocation. The COPY instruction is an assignment statement. PHI is the convergence point of a branching structure that selects different values based on the branch taken. CAST is used to interpret an object as a specific type to satisfy LLVM’s type checking without altering the contents in memory. LOAD and STORE are operations for reading from or writing to address-taken objects. FIELD obtains the address of an object’s field. CALL and RET instructions are used for function calls and returns.
The symbols used in this paper for pointer analysis are as follows: $pt(p)$ represents the set of objects pointed to by pointer $p$. For example, $pt(p)=\{o1,o2\}$ means that pointer $p$ points to objects $o1$ and $o2$. $pt.t(o)$ represents the set of types owned by object $o$. $pt.o(t)$ represents the set of field offsets that type $t$ owns. The function $off()$ is used to obtain the offset of a field within an object. The symbol $\widehat{o_{f}}$ denotes the sub-object of object $o$ with an offset of $f$. The symbol * represents any field offset. The rest of the symbols have the same meaning as mathematical set operators.
4.1.2. Pointer Analysis
The instructions involved in pointer analysis can be categorized into five types: object creation instructions, object propagation instructions, type conversion instructions, field access instructions, and function call and return instructions. The inference rules for the first phase of pointer analysis are shown in Figure 4.
@@ 4.1.2. Pointer Analysis
$\begin{array}[]{l}\textbf{{\scriptsize[STACK/GLOBAL]}}\\ $$\ \ p=alloca\ t,nbytes$$\\ \hline\cr$$\ \ pt(p)=pt(p)\cup\{o\}\\ pt.t(o)=pt.t(o)\cup\{t\}$$\end{array}$ $\quad\ \ \ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[HEAP]}}\\ $$\ \ p=malloc\ nbytes$$\\ \hline\cr pt(p)=pt(p)\cup\{o\}$$\end{array}$ $\quad\ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[COPY]}}\\ $$\ \ \ \ \ \ p=q$$\\ \hline\cr$$pt(q)\subseteq pt(p)$$\end{array}$ $\quad\ \ \ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[CAST]}}\\ $$p=(t)q\ \ \ \ \ o\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ pt(q)\subseteq pt(p)\\ pt.t(o)=pt.t(o)\cup\{t\}$$\end{array}$
@@
$\begin{array}[]{l}\textbf{{\scriptsize[PHI]}}\\ $$\ \ \ \ \ \ \ \ p=\phi(q,x)$$\\ \hline\cr$$pt(q)\cup pt(x)\subseteq pt(p)$$\end{array}$ $\quad\ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[LOAD]}}\\ $$p=*q\ \ o\in pt(q)$$\\ \hline\cr$$\ \ pt(o)\subseteq pt(p)$$\end{array}$ $\quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[STORE]}}\\ $$*p=q\ \ o\in pt(p)$$\\ \hline\cr$$\ \ pt(q)\subseteq pt(o)$$\end{array}$ $\quad\ \ \quad\ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[FIELD-VAR]}}\\ $$p=\&q\rightarrow*\ \ \ o\in pt(q)$$\\ \hline\cr$$\ \ pt(p)=pt(p)\cup\{\widehat{o_{\ast}}\}$$\end{array}$
@@
$\begin{array}[]{l}\textbf{{\scriptsize[FIELD]}}\\ $$p=\&q\rightarrow fld\ \ \ f=off(q\rightarrow fld)\ \ \ o\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ \ \ \ \ \ \ \ \ \ pt(p)=pt(p)\cup\{\widehat{o_{f}}\}$$\end{array}$ $\quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[CALL]}}\\ $$p=q(...,a,...)\ \ m(...,a^{\prime}...)\ \ m\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ \ \ \ \ \ \ \ pt(a)\subseteq pt(a^{\prime})$$\\ \end{array}$ $\begin{array}[]{l}\textbf{{\scriptsize[RET]}}\\ $$p=q(...)\ \ \ ret\ p^{\prime}\ \ \ m\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ \ \ pt(p^{\prime})\subseteq pt(p)$$\end{array}$
Figure 4. Stage 1 Pointer Analysis Rules.
@@
Object creation involves two rules: [STACK/GLOBAL] and [HEAP]. [STACK/GLOBAL] is used to create stack and global objects, and it can obtain the type information of objects from the creation instruction. In this rule, an object $o$ is created, and we use $o\in pt(p)$ to indicate that pointer $p$ points to object $o$. The type of object $o$ in this rule is $t$, and we use $t\in pt.t(o)$ to indicate that the object has type $t$. The [HEAP] instruction only involves the creation of objects.
@@
Object propagation involves the four rules [COPY], [PHI], [LOAD], and [STORE]. These rules only involve the propagation of objects. Both the [COPY] and [PHI] rules are simple assignment instructions, propagating the points-to information of the right-hand pointer to the left side. Specifically, the [PHI] rule merges the points-to sets of the two pointers on the right and propagates them to the left. [LOAD] and [STORE] propagate objects with the help of a third pointer.
@@
Function calls involve the [CALL] and [RET] rules. We employs an On-The-Fly approach for resolving function calls. The On-The-Fly further analyzes the call relationships between functions based on the results of pointer analysis. This kind of function call analysis can effectively analyze indirect function calls, such as $call$ $*q$.
$\begin{array}[]{l}\textbf{{\scriptsize[FIELD]}}\\ $$\ \ \ \ \ \ \ \ p=\&q\rightarrow fld\ \ \ \ \ \ \ \ \ \ f=off(q\rightarrow fld)$$\\ \hline\cr$$\forall o_{i}\in pt(q)\land t\in pt.t(o_{i})\land f\in pt.o(t),pt(p)=pt(p)\cup\{\widehat{o_{f}}\}$$\\ \end{array}$
Figure 5. Stage 1 Pointer Analysis Rules.
@@
The specific pointer analysis process involves the following steps: During stage 1 of pointer analysis, we repeatedly apply the pointer analysis rules in Figure xxx until the pointer pointing set no longer changes, marking the end of stage 1 pointer analysis. Then, stage 2 of pointer analysis begins, which is similar to stage 1 but uses the object type information obtained from the first stage. The pointer analysis concludes when the set of pointer pointing set no longer changes.
@@
4.2. Mapping Static Results to Runtime
In this section, we explain the method of mapping static analysis results to runtime information of the program. First, we describe the method of mapping static objects to runtime objects. Then, we explain the method of mapping static pointers to runtime pointers. Finally, we explain our method for accurately eliminating dangling pointers by utilizing the pointer analysis results and the program’s runtime information.
4.2.1. Object Mapping
There are two reasons why we need to map static objects to runtime objects: 1. A single static object corresponds to multiple runtime objects. 2. The memory addresses of objects can only be determined at runtime. During pointer analysis, we abstract each heap allocation point as a specific object. This abstraction method allows us to quickly analyze points-to relationships during pointer analysis, but it is not precise enough for eliminating dangling pointers. This is because the same allocation point can create different objects in different contexts. Therefore, it is unreasonable to continue using this abstraction method when eliminating dangling pointers.
To distinguish different objects created at the same allocation point, we use the runtime information of the program. We insert a function after each allocation point in the program. The main function of this inserted function is to store the allocated memory information in our designed multithreaded hash table and to pass the information of the created objects to our thread through a multithreaded lock-free queue (Aldinucci et al., 2012). The LightDE thread groups all dynamic objects created at the same allocation point together, enabling us to find the corresponding runtime objects based on the static object. The object information stored in our data structure includes the memory start address, end address, size, and the identifier of the object’s corresponding static object.
4.2.2. Pointer Mapping
@@
Heap Pointer Mapping Before introducing our method for mapping heap pointers, we use an example to illustrate the relationship between static heap pointers and runtime heap pointers. Suppose there are two objects, $o1$ and $o2$. The $f$ field of $o2$ points to the object $o1$. In static analysis, $o1$ and $o2$ correspond to two allocation points, which may correspond to many runtime objects. Our mapping method involves allocating a data structure for each static object based on the pointer analysis results during program compilation. The data structure records the heap pointers pointing to the object. Since the addresses of heap pointers cannot be determined during compilation, we record the static identifier of the object containing the heap pointer and the offset of the heap pointer within the static object in the data structure. When an object is released, our method for eliminating heap pointers that point to the released object is to check the heap pointers at the same offset location in the runtime objects corresponding to the static object containing the heap pointer. If we cannot determine the exact offset of the heap pointer within the object, we check all the object’s pointer fields to see if they point to the freed object. Since our pointer analysis is structure-sensitive, we can determine all the offsets of the object’s pointer fields.
@@
Global Pointer Mapping In this paragraph, we explain our method for mapping static global pointers to runtime global pointers. Global pointers can be accessed from any location in the program, but their memory addresses cannot be determined during program compilation. To address this issue, we assign indices based on the names of global pointers. During compilation, we store the indices of global pointers that point to a certain static object in its pointer set. We insert a function at the beginning of the program’s startup function, which stores the memory addresses of global pointers into a global array according to their indices. Using this global array, we can check whether the global pointers identified by pointer analysis point to released runtime objects based on the objects’ static identifiers.
@@
Stack Pointer Mapping In this paragraph, we explain our method for mapping static stack pointers to runtime stack pointers. Our method involves assigning a static identifier to each function in the program during compilation, and assigning a static identifier to all stack pointers within each function in the order they are defined. During compilation, we store the function identifier of the stack pointer, as well as the stack pointer’s static identifier within the function, into the object’s pointer set. In each function, we insert instructions to store the stack pointer into our data structure based on the static identifier within the function. When an object is released, we check whether the stack pointers identified by the pointer analysis point to the released object. Since we use a separate thread to asynchronously eliminate stack pointers, we need to avoid the stack space being freed while eliminating stack pointers. We designed a dedicated stack for this purpose. When the program needs to allocate stack pointers, it allocates space on this dedicated stack. However, the original thread of the program does not release the allocated space. When a function exits, the program’s original thread generates exit information and passes it to our thread. Our thread then releases the corresponding stack space.
4.2.3. Dangling Pointer Elimination
@@
In this paragraph, we explain the method LightDE uses to eliminate corresponding dangling pointers after an object is released. We insert our memory management function before the release points identified by the pointer analysis and remove the program’s original memory release function. The LightDE thread releases the program’s allocated memory to prevent UAF vulnerabilities from occurring during the elimination of dangling pointers by LightDE. Our memory management function removes the information of the released object from our multithreaded hash table and passes this information to the LightDE thread. The LightDE thread then retrieves the static identifier of this runtime object and uses it to obtain the static pointers stored in the program’s data segment that point to the object. Next, the LightDE thread uses the pointer mapping method described in 4.2.2 to check whether the corresponding runtime pointers point to the released object, and it invalidates the corresponding dangling pointers.
5. Evaluation

### v4

@@ IV-A2 Program Abstraction
TABLE I: Analysis Domains.
@@
$m$in $\mathcal{M}$ program functions
@@
The domains related to our pointer analysis are shown in Table I, mainly including variables, functions, types, and objects. Variables in programs $\mathcal{V}$ can be divided into two types: address-taken variables $\mathcal{A}$, which can be accessed through pointers, and top-level variables $\mathcal{P}$, which are never accessed through pointers. Address-taken variables $\mathcal{A}$ include abstract objects $\mathcal{O}$ and fields of abstract objects $\mathcal{Y}$. Heap objects are a type of address-taken object, and our method for abstracting heap objects is to treat each heap allocation point as a distinct object. We represent object fields by creating a sub-object for each field and establishing a relationship between the sub-object and its parent object. The creation of sub-objects is based on the offset of their fields relative to the base object.
TABLE II: LLVM-like Instructions.
@@
$\begin{array}[]{l}\textbf{{\scriptsize[STACK/GLOBAL]}}\\ $$\ \ p=alloca\ t,nbytes$$\\ \hline\cr$$\ \ pt(p)=pt(p)\cup\{o\}\\ pt.t(o)=pt.t(o)\cup\{t\}$$\end{array}$ $\quad\ \ \ \ \ \ \begin{array}[]{l}\textbf{{\scriptsize[HEAP]}}\\ $$\ \ p=malloc\ nbytes$$\\ \hline\cr pt(p)=pt(p)\cup\{o\}$$\end{array}$ $\quad\ \ \ \ \begin{array}[]{l}\textbf{{\scriptsize[COPY]}}\\ $$\ \ \ \ \ \ p=q$$\\ \hline\cr$$pt(p)\supseteq pt(q)$$\end{array}$ $\quad\ \ \ \ \ \begin{array}[]{l}\textbf{{\scriptsize[CAST]}}\\ $$p=(t)q\ \ \ \ \ o\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ pt(p)\supseteq pt(q)\\ pt.t(o)=pt.t(o)\cup\{t\}$$\end{array}$
@@
$\begin{array}[]{l}\textbf{{\scriptsize[PHI]}}\\ $$\ \ \ \ \ \ p=\phi(q,x)$$\\ \hline\cr$$pt(p)\supseteq pt(q)\cup pt(x)$$\end{array}$ $\quad\ \ \ \ \begin{array}[]{l}\textbf{{\scriptsize[LOAD]}}\\ $$p=*q\ \ o\in pt(q)$$\\ \hline\cr$$\ \ pt(p)\supseteq pt(o)$$\end{array}$ $\quad\ \ \ \begin{array}[]{l}\textbf{{\scriptsize[STORE]}}\\ $$*p=q\ \ o\in pt(p)$$\\ \hline\cr$$\ \ pt(o)\supseteq pt(q)$$\end{array}$ $\quad\ \ \quad\ \ \ \begin{array}[]{l}\textbf{{\scriptsize[FIELD-VAR]}}\\ $$p=\&q\rightarrow*\ \ \ o\in pt(q)$$\\ \hline\cr$$\ \ pt(p)=pt(p)\cup\{\widehat{o_{\ast}}\}$$\end{array}$
@@
$\begin{array}[]{l}\textbf{{\scriptsize[FIELD]}}\\ $$p=\&q\rightarrow fld\ \ f=off(q\rightarrow fld)\ \ o\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ \ \ \ \ \ \ \ pt(p)=pt(p)\cup\{\widehat{o_{f}}\}$$\end{array}$ $\ \ \begin{array}[]{l}\textbf{{\scriptsize[CALL]}}\\ $$p=q(...,a,...)\ \ m(...,a^{\prime},...)\ \ m\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ \ \ \ \ \ \ \ pt(a^{\prime})\supseteq pt(a)$$\\ \end{array}$ $\begin{array}[]{l}\textbf{{\scriptsize[RET]}}\\ $$p=q(...)\ \ ret\ p^{\prime}$$\\ \hline\cr$$\ \ pt(p)\supseteq pt(p^{\prime})$$\end{array}$
Figure 6: Stage 1 Pointer Analysis Rules.
Our pointer analysis is based on the LLVM IR of programs. LLVM defines a rich set of instructions, but we can abstract these instructions into a small instruction set to represent all program instructions. Our abstracted instruction set is shown in Table II. STACK/GLOBAL instructions allocate stack or global objects. HEAP instructions are used for heap object allocation. The COPY instruction is an assignment statement. PHI is the convergence point of a branching structure that selects different values based on the branch taken. CAST is used to interpret an object as a specific type to satisfy LLVM’s type checking without altering the contents in memory. LOAD and STORE are operations for reading from or writing to address-taken objects. FIELD obtains the address of an object’s field. CALL and RET instructions are used for function calls and returns.
The symbols used in this paper for pointer analysis are as follows: $pt(p)$ represents the set of objects pointed to by pointer $p$. For example, $pt(p)=\{o1,o2\}$ means that pointer $p$ points to objects $o1$ and $o2$. $pt.t(o)$ represents the set of types owned by object $o$. $pt.o(t)$ represents the set of field offsets that type $t$ owns. The function $off()$ is used to obtain the offset of a field within an object. The symbol $\widehat{o_{f}}$ denotes the sub-object of object $o$ with an offset of $f$. The symbol * represents any field offset of an object. The symbol $\supseteq$ denotes a containment relationship, meaning the elements of the right set are added to the left set. For example, $pt(p)\supseteq pt(q)$ means the elements of the set $pt(q)$ are added to the set $pt(p)$. The rest of the symbols have the same meaning as the mathematical set operators.
IV-A3 Pointer Analysis
The instructions for pointer analysis can be categorized into five types: object creation, object propagation, type conversion, field access, and function call/return instructions. The inference rules for the first phase of pointer analysis are shown in Figure 6.
@@ IV-A3 Pointer Analysis
Object creation involves two rules: [STACK/GLOBAL] and [HEAP]. [STACK/GLOBAL] is used to create stack and global objects, and it can obtain the type information of objects from the creation instruction. In this rule, an object $o$ is created, and we use $pt(p)=pt(p)\cup\{o\}$ to indicate that the pointer $p$ points to a new object $o$. The type of object $o$ in this rule is $t$, and we use $pt.t(o)=pt.t(o)\cup\{t\}$ to indicate that the object acquires type $t$. The [HEAP] rule only involves the creation of heap objects.
@@
Object propagation involves the four rules [COPY], [PHI], [LOAD], and [STORE]. These rules only involve the propagation of objects. Both the [COPY] and [PHI] rules are simple assignment instructions, propagating the points-to information of the right-hand pointer to the left-hand pointer. Specifically, the [PHI] rule merges the points-to sets of the two pointers on the right and propagates them to the left-hand pointer. [LOAD] and [STORE] propagate objects with the help of a third pointer.
@@
Function calls involve the [CALL] and [RET] rules. We employ an On-The-Fly approach for resolving function calls. This method further analyzes the call relationships between functions based on the results of pointer analysis. This kind of function call analysis can effectively analyze indirect function calls, such as $call$ $*q$.
$\begin{array}[]{l}\textbf{{\scriptsize[FIELD]}}\\ $$\ \ \ \ \ \ \ \ p=\&q\rightarrow fld\ \ \ \ \ \ \ \ \ \ f=off(q\rightarrow fld)$$\\ \hline\cr$$\forall o_{i}\in pt(q)\land t\in pt.t(o_{i})\land f\in pt.o(t),pt(p)\supseteq\{\widehat{o_{if}}\}$$\\ \end{array}$
Figure 7: Stage 2 Pointer Analysis Rule.
@@
The specific pointer analysis process involves the following steps: During stage 1 of pointer analysis, we repeatedly apply the pointer analysis rules in Figure 6 until the pointer pointing set no longer changes, marking the end of stage 1. Then, stage 2 of pointer analysis begins. This stage is similar to stage 1 but uses the object type information obtained from the first stage to make the pointer analysis more accurate. The pointer analysis concludes when the pointer pointing set no longer changes.
@@
IV-B Mapping Static Results to Runtime
In this section, we explain the method of mapping static pointer analysis results to runtime information of the program. First, we describe the method of mapping static objects to runtime objects. Then, we explain the method of mapping static pointers to runtime pointers. Finally, we explain our method for accurately eliminating dangling pointers by utilizing the pointer analysis results and the program’s runtime information.
IV-B1 Object Mapping
There are two reasons why we need to map static objects to runtime objects: 1. A single static object corresponds to multiple runtime objects. 2. The memory addresses of objects can only be determined at runtime. During pointer analysis, we abstract each heap allocation point as a specific object. This abstraction method allows us to quickly analyze points-to relationships during pointer analysis, but it is not precise enough for eliminating dangling pointers. This is because the same allocation point can create multiple different objects based on different contexts. Therefore, it is unreasonable to continue using this abstraction method when eliminating dangling pointers.
To distinguish different objects created at the same allocation point, we use the runtime information of the program. We insert a function after each allocation point in the program. The main purpose of this inserted function is to store the allocated memory information in our designed multi-threaded hash table and to pass the information of the created objects to our thread through a multi-threaded lock-free queue [38, 39]. The LightDE thread groups all dynamic objects created at the same allocation point together, enabling us to find the corresponding runtime objects based on the static object. The object information stored in our data structure includes the memory start address, end address, size, and the identifier of the corresponding static object.
IV-B2 Pointer Mapping
@@
Heap Pointer Mapping Before introducing our method for mapping heap pointers, we use an example to illustrate the relationship between static heap pointers and runtime heap pointers. Suppose there are two objects, $o1$ and $o2$. The $f$ field of $o2$ points to the object $o1$. In static analysis, $o1$ and $o2$ correspond to two allocation points, which may correspond to many runtime objects. Therefore, $o2.f$ corresponds to multiple heap pointers at runtime. Our mapping method involves allocating a data structure for each static object based on the pointer analysis results during program compilation. The data structure records the heap pointers pointing to the object. Since the addresses of heap pointers cannot be determined during compilation, we record the static identifier of the static object containing the heap pointer and the offset of the heap pointer within the static object in the data structure. When an object is released, our method for eliminating heap pointers that point to the released object is to check the heap pointers at the same offset location in the runtime objects corresponding to the static object that contains the heap pointers. If we cannot determine the exact offset of the heap pointer within the object, we check all the object’s pointer fields to see if they point to the freed object. Since our pointer analysis is structure-sensitive, we can determine all the offsets of the object’s pointer fields.
@@
Global Pointer Mapping In this paragraph, we explain our method for mapping static global pointers to runtime global pointers. Global pointers can be accessed from any location in the program, but their memory addresses cannot be determined during program compilation. To address this issue, we assign indices based on the names of global pointers. During compilation, we store the indices of global pointers that point to a certain static object in its pointer set. We insert a function at the beginning of the program’s startup function, which stores the memory addresses of global pointers into a global array according to their indices. Using this global array, we can check whether the global pointers that point to a static object are pointing to the corresponding released runtime object.
@@
Stack Pointer Mapping In this paragraph, we explain our method for mapping static stack pointers to runtime stack pointers. Our method involves assigning a static identifier to each function of the program during compilation and assigning a static index to all stack pointers within each function in the order they are defined. During compilation, we store the function identifier that the stack pointer resides in and the stack pointer’s static index within the function into the object’s pointer set based on the pointer analysis results. In each function, we insert instructions to store the stack pointer into our data structure based on the static index within the function. When an object is released, we check whether the stack pointers identified by the pointer analysis point to the released object. Here, we can also see the shortcomings of our method for handling stack pointers. Our method still requires storing the memory addresses of stack pointers at runtime, but it avoids looking up the objects pointed to by the pointers. Since we use a separate thread to asynchronously eliminate stack pointers, we need to avoid the stack space being freed while eliminating stack pointers. We designed a dedicated stack for this purpose. When the program needs to allocate stack pointers, it allocates space on this dedicated stack. However, the original thread of the program does not release the allocated space. When a function exits, the program’s original thread generates exit information and passes it to our thread. Our thread then releases the corresponding stack space.
IV-B3 Dangling Pointer Elimination
@@
In this paragraph, we explain the method LightDE uses to eliminate corresponding dangling pointers after an object is released. We insert our memory management function before the release points identified by the pointer analysis and remove the program’s original memory release function. The LightDE thread releases the memory after setting the pointers pointing to the released memory to invalid pointers to prevent UAF vulnerabilities during the elimination of dangling pointers. Our memory management function removes the information of the released object from our multi-threaded hash table and passes this information to the LightDE thread. The LightDE thread then retrieves the static identifier of this runtime object and uses it to obtain the static pointers stored in the program’s data segment that point to the object. Next, the LightDE thread uses the pointer mapping method described in IV-B2 to check whether the corresponding runtime pointers point to the released object, and it invalidates the corresponding dangling pointers.
TABLE III: Comparison the Accuracy of Our Structure-sensitivity with Andersen’s Pointer Analysis.
Benchmark
Andersen
Points-to#
Our method
Points-to#
Points-to#
Filtered out
400.perlbench 82704268 82195642 508626
401.bzip2 23845 23097 748
403.gcc 516822633 512698520 4124113
429.mcf 2293 2071 222
433.milc 30626 30454 172
444.namd 21422 20347 1075
445.gobmk 518615 518588 27
447.dealII 9557789 9033265 524524
450.soplex 12375851 11314009 1061842
453.povray 45313149 39046254 6266895
456.hmmer 218420 208138 10282
458.sjeng 12398 12281 117
462.libquantum 2975 2948 27
464.h264ref 1484659 1464771 19888
470.lbm 789 789 0
471.omnetpp 154075153 71838897 82236256
473.astar 5025 5008 17
482.sphinx3 227546 209173 18373
483.xalancbmk 126216408 119172228 7044180
V Evaluation